Can someone please help me:
Solve the inequality.
4x/6-x (less than) x

Answer by stanbon(75887)

Solve the inequality.
4x/6-x (less than) x
--------------------------
Solve the equality:
4x/(6-x) = x
4x = x(6-x)
4x = 6x - x^2
x^2 - 2x = 0
x(x-2) = 0
x = 0 or x = 2
Comment x = 6, x = 0, and x = 2 are values that x cannot take.
----------------------------
Draw a number line; plot the numbers 0,2,6 on the line.
Pick a test value in each of the 4 line intervals to see
where the solutions lie.
-----------------------------------
4x/6-x (less than) x
-----------------------------------
Check x = -1; -4/7 < -1; false
Check x = 1 ; 4/(5)<1 ; true
Check x = 4 ; 16/2<4 ; false
Check x = 8 ; 21/2 < 8 ; false
-----------------------------------
Solution:
0 < x < 2
===========================
Cheers,
Stan H.